OpenStudy (anonymous):

Find the lateral area of a cone with a radius of 7 ft. and a slant height of 13 ft. Use 3.14 for π and round to the nearest tenth. 439.6 ft2 324.5 ft2 571.5 ft2 285.7 ft2

OpenStudy (mertsj):

What is the formula for the lateral area of a cone?

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

@ mertsj it's v=1/3 bh

OpenStudy (mertsj):

That is the formula for the volume. Notice your problem asks for lateral area. It's a different formula.

OpenStudy (anonymous):

L.a. =1/2 (2pi r)l = pi r l

OpenStudy (mertsj):

That's it!!! What is r in this problem?

OpenStudy (anonymous):

Radius is 7 ft.

OpenStudy (mertsj):

What is l ?

OpenStudy (anonymous):

l is the height of the cone

OpenStudy (mertsj):

No. It is the slant height.

OpenStudy (anonymous):

Oh so 13ft.

OpenStudy (mertsj):